A demolition company wants to build a brick wall to hide from public view of the area where they store wrecked cars. Working together, an experienced bricklayer and an apprentice can build the wall in 12 hours. Working alone, it takes the apprentice 10 hours longer than the experienced bricklayer to do the job. How long would it take the experienced bricklayer to build the wall working alone?
Answer by checkley77(12844)

1/12=1/x+1/(x+10)
1/12=(x+10+x)/x(x+10)
1/12=(2x+10)/(x^2+10x)
12(2x+10)=x^2+10x
24x+120=x^2+10x
x^2+10x-24x-120=0
x^2-14x-120=0
(x-20)(x+6)=9
x-20=0
x=20 hours for the experienced bricklayer.
Proof:
1/12=1/20+1/30
1/12=(30+20)/20*30
1/12=50/600
12*50=600
600=600